OpenStudy (anonymous):

Write the equation of the sine function with a period of pi, an amplitude of 2 , a right phase shift of pi/2, and vertical translation up 1 unit.

OpenStudy (calculusfunctions):

Use the fact that\[y =a \sin \sin [k(x -d)]+c\]where the parameters a, d, and c represent amplitude, phase shift, and vertical translation respectively and the parameter\[k =\frac{ 2\pi }{ period }\]